Mrs. Garcia is 1 year less than 3 times as old as her daughter . Seven years from now, she will be 4 years more than twice as old as her daughter will be then. Find both of their ages.
Answer by scott8148(6628)

m = 3d - 1

m + 7 = 2(d + 7) + 4

substituting ___ (3d - 1) + 7 = 2(d + 7) + 4 ___ 3d + 6 = 2d + 18